# Gravitational odd-parity perturbation of a Horndeski hairy black hole: quasinormal mode and parameter constraint

### v2

@@ II.2 Odd-parity perturbation and modified Regge-Wheeler equation
in which we have integrated out the coordinates $\theta$ and $\phi$. Here the dot and prime denote the derivative with respect to $t$ and $r$, respectively, and the coefficients are expressed as
$a_{1}$ $=$ $\frac{\ell(\ell+1)}{4r^{2}}\left[\left[4r(Q_{4}-2\chi Q_{4,\chi})\right]^{\prime}+\frac{2(\ell-1)(\ell+2)}{f}Q_{4}\right],$ (14)
$a_{2}$ $=$ $-\frac{\ell(\ell+1)(\ell-1)(\ell+2)f}{2r^{2}}(Q_{4}-2\chi Q_{4,\chi}),$ (15)
$a_{3}$ $=$ $\frac{\ell(\ell+1)}{2}(Q_{4}-2\chi Q_{4,\chi}).$ (16)
To proceed, we introduce a new auxiliary field De Felice et al. (2011); Ganguly et al. (2018)
$\Psi(r,t)=\frac{\sqrt{2}r(Q_{4}-2\chi Q_{4,\chi})}{\sqrt{Q_{4}}}\left(\dot{h}_{1}-h^{\prime}_{0}+\frac{2h_{0}}{r}\right)$ (17)
such that under the tortoise coordinate $r_{*}=\int\frac{dr}{f}$, the second order action (13) can be reduced as
[…]
where the potential $V(r)$ takes the form
$V(r)=\ell(\ell+1)\frac{fQ_{4}}{r^{2}(Q_{4}-2\chi Q_{4,\chi})}-\frac{r^{3}}{2}\left(\frac{f^{2}}{r^{4}}\right)^{\prime}-\frac{r^{4}Q_{4}^{2}}{4Q_{4}^{\prime}}\left(\frac{f^{2}Q_{4}^{\prime 2}}{r^{4}Q_{4}^{3}}\right)^{\prime}-\frac{2fQ_{4}}{r^{2}(Q_{4}-2\chi Q_{4,\chi})}.$ (19)
It is obvious from (18) that the second order actions is only well-defined for $\ell\geq 2$ but not for the modes $\ell=0$ and $\ell=1$ which are in fact not dynamical.
Taking the variation of the action (18), we can obtain the equation of motion
$\left(\frac{d}{dr_{*}^{2}}-\frac{Q_{4}}{Q_{4}-2\chi Q_{4,\chi}}\frac{d}{dt^{2}}-V(r)\right)\Psi=0.$ (20)
which is the modified Regge-Wheeler equation for the Horndeski hairy black hole (11) in the subclass of Horndeski theory with the action (2). Solving this equation with proper boundary conditions allows us to extract the spectrum of the odd-parity perturbation and to further analyze the stability of the hairy black hole and other related properties, which will be discussed in the next sections. Before closing this section, we will briefly examine the radial speed of gravitational wave in the current model, which is also known as the speed of perturbation propagating along the the radial direction. This speed is defined by the coefficient of $\frac{d}{dt^{2}}$ as
$c_{r}^{2}=\frac{Q_{4}-2\chi Q_{4,\chi}}{Q_{4}}=\left(1+\frac{3Q}{2r}\right)^{-1}.$ (21)
@@
It is obvious that the scalar charge $Q$ indeed cause the deviation of $c_{r}^{2}$ from the vacuum light speed, i.e. the result in GR. Equation (21) shows that the Laplacian instability caused by $c_{r}^{2}<0$ may occur. Considering that it is a monotonically increasing function of the radial coordinate $r$, so in order to avoid such instability, a condition for keeping a positive $c_{r}^{2}$ must be imposed at least at the event horizon, i.e.
@@
which is consistent with the stability condition addressed in Kobayashi et al. (2012); Ganguly et al. (2018). Moreover, to illustrate how the hairy parameter affect the radial propagation speed of the odd perturbation, we plot FIG.1 which shows that propagation speed can be larger or smaller than the vacuum light speed depending on the hairy charge.
@@ III Quasinormal mode
Recalling the Horndeski hairy black hole (11), the modified Regge-Wheeler equation can be rewritten as
$\left[\frac{d}{dr_{*}^{2}}-\left(1+\frac{3Q}{2r}\right)\frac{d}{dt^{2}}-V(r)\right]\Psi=0~{}~{}~{}~{}\text{with}~{}~{}~{}~{}$ (24)
$V(r)=\frac{f^{2}\left(27Q^{2}+48Qr+32r^{2}\right)}{4r^{2}(3Q+2r)^{2}}+\left(1+\frac{3Q}{2r}\right)f\left[\frac{(\ell-2)(\ell+1)}{r^{2}}-\frac{f^{\prime}(3Q+4r)}{(3Q+2r)^{2}}\right],$
which returns to the Regge-Wheeler potential Regge and Wheeler (1957) when the Horndeski hair vanishes. Since the explicit expression for the wave-like equation is clearly established, we are now in a position to solve the equation and extract the dynamical properties for the odd-parity perturbation, which allows us to work in both time and frequency domains.
@@ III.1 Preparation
Firstly, we will give the method and boundary conditions in our numerical calculations to solve (24) from two aspects.
@@
  • In time domain, we shall work in tortoise coordinate grids and apply the finite difference method (the readers may refer to Zhu et al. (2014); Yang et al. (2024) for more details) to discretize the radial coordinate $r=r(r_{*})=r(j\Delta r_{*})=r_{j}$, thus the perturbation and the potential are discretized as $\Psi(t,r)=\Psi(i\Delta t,j\Delta r_{*})=\psi_{i,j}$ and $V(r)=V(j\Delta r_{*})=V_{j}$. Then we use the central difference to discretize the derivative of the perturbation and rewrite the equation (24) into an iterative equation
    $\Psi_{i+1,j}=\frac{\Delta t^{2}\left(\Psi_{i,j-1}-2\Psi_{i,j}+\Psi_{i,j+1}\right)}{\left(1+\frac{3Q}{2r_{j}}\right)\Delta r_{*}^{2}}-\Delta t^{2}V_{j}\Psi_{i,j}+2\Psi_{i,j}-\Psi_{i-1,j}.$ (25)
    By imposing the initial conditions $\Psi_{i<0,j}=0$, $\Psi_{0,j}=exp[-\frac{(r_{j}-18)^{2}}{2}]$, and setting $\frac{\Delta t}{\Delta r_{*}}=\frac{0.05}{0.1}=\frac{1}{2}$ to satisfy the Von Neumann stability, we can finally solve the evolution of the odd perturbation from the iterative equation. The evolutionary profile of the perturbation typically includes an initial outburst, a damped oscillation, and a late-time tail, from which one can compare the damping time of the perturbation in varying black hole parameters and roughly predict the (in)stability of the background spacetime.
@@
  • In frequency domain, we shall factor out the time dependency of the perturbation by separating $\Psi(t,r)=e^{-i\omega t}R(r)$ in (24), which yields a general eigenvalue equation
@@
    $\left[\frac{d}{dr_{*}^{2}}+\left(1+\frac{3Q}{2r}\right)\omega^{2}-V(r)\right]R(r)=0.$ (26)
@@
    The boundary conditions for the QNMs solution are ingoing at the horizon $R(r\to r_{+})\sim e^{-i\sqrt{1+\frac{3Q}{2r_{+}}}\omega r_{*}}$ and outgoing at the infinity $R(r\to\infty)\sim e^{i\omega r_{*}}$, respectively. Thus, once the black hole parameters (for instance, the mass $M$ and the hairy charge $Q$) are fixed, the eigenvalues could be solved as an infinite series of complex numbers, which are the so-called QNM frequencies. Similar to the normal modes, the real part of the QNMs represents the physical oscillation angular frequency, but they differ from the normal modes by having a non-vanishing imaginary part that represents the exponential damping of the perturbation in a dissipating system.
@@
    We shall employ the pseudo-spectral method and the matrix method to solve the master equation (26) and determine the QNMs for the gravitational odd modes of the Horndeski black hole. Since both algorithms are widely used in studying of black hole spectroscopy, here we will not review the main steps, and readers may refer to Jansen (2017) for the pseudo-spectral method and Lin and Qian (2019) for the matrix method, respectively. It is noticed that due to the $log$ term in the metric function, a redefinition of the radial wave function is crucial when using both methods, as we have addressed in appendix B of our previous work Yang et al. (2024). Similar processes will be done in the current numerical calculations.
@@
Following the above preparation, we shall show our numerical results of the eigenfrequencies of the QNMs and time evolution of perturbed field in next subsection. In our numeric, we find that as the hairy parameter goes closer to the extremal case, the precision of our calculation becomes worse and even the code may break down. So, we shall focus on the parameter regime $-M\leq Q\leq M$, and for simplicity, we also rescale the Horndeski charge into the dimensionless $\tilde{Q}=Q/M$ in the calculations.
@@ III.2 Numerical results
The QNM frequencies with nodes $n=0,~{}1$ for the selected hairy charges obtained from the pseudo-spectral method and the matrix method are listed in TABLE 1, where we also pledge the accuracy of the results by evaluating their relative error. It is shown that for each mode, when $|\tilde{Q}|$ is small, both methods achieve high accuracy in the nodes $n=0~{}~{}and~{}~{}1$, and the results are in good agreement with each other at least at the order of $10^{-6}$ or even better.
@@
$\ell=2$ $n=0$ 0.747343 - 0.177925 i 0.679390 - 0.163833 i 0.617493 - 0.153284 i
0.747343 - 0.177925 i 0.679390 - 0.163833 i 0.617495 - 0.153284 i
[…]
$n=1$ 0.693422 - 0.547830 i 0.623772 - 0.504366 i 0.573986 - 0.481454 i
0.693420 - 0.547829 i 0.623768 - 0.504368 i 0.574023 - 0.481421 i
[…]
$\ell=3$ $n=0$ 1.198887 - 0.185406 i 1.078947 - 0.167580 i 0.955588 - 0.150350 i
1.198887 - 0.185406 i 1.078947 - 0.167580 i 0.955587 - 0.150349 i
[…]
$n=1$ 1.165288 - 0.562596 i 1.046267 - 0.508376 i 0.929446 - 0.458644 i
1.165288 - 0.562596 i 1.046269 - 0.508377 i 0.929442 - 0.458645 i
[…]
$\ell=4$ $n=0$ 1.618357 - 0.188328 i 1.451873 - 0.169314 i 1.274651 - 0.149665 i
1.618357 - 0.188328 i 1.451873 - 0.169314 i 1.274651 - 0.149666 i
[…]
$n=1$ 1.593263 - 0.568668 i 1.427902 - 0.511200 i 1.255580 - 0.452830 i
1.593263 - 0.568669 i 1.427904 - 0.511200 i 1.255583 - 0.452831 i
[…]
Table 1: The QNMs obtained by pseudo-spectral (upper) and matrix (lower) method of the hairy black hole, for $\ell=2,3,4$ and $n=0,1$. The relative errors are also shown below each mode, defined by $\Delta_{k}=\frac{\text{Matrix}(\omega_{k})-\text{Pseudospectral}(\omega_{k})}{\text{Pseudospectral}(\omega_{k})}100\%$ with $k=\text{Re},\text{Im}$.

### v3

@@ II.2 Odd-parity perturbation and modified Regge-Wheeler equation
in which we have integrated the angular distributions with respect to $\theta$ and $\phi$. Here the dot and prime denote the derivative with respect to $t$ and $r$, respectively, and the coefficients are expressed as
$a_{1}$ $=$ $\frac{\ell(\ell+1)}{4C}\left[\partial_{r}\left(C^{\prime}\sqrt{\frac{B}{A}}{\cal{H}}\right)+\frac{(\ell-1)(\ell+2)}{\sqrt{AB}}\cal{F}\right],$ (14)
$a_{2}$ $=$ $-\frac{\ell(\ell+1)(\ell-1)(\ell+2)\sqrt{AB}}{2C}\cal{G},$ (15)
$a_{3}$ $=$ $\frac{\ell(\ell+1)}{4}\sqrt{\frac{B}{A}}{\cal{H}},\quad\text{with}$ (16)
${\cal F}$ $=$ $2\left(Q_{4}+\frac{1}{2}B\phi^{\prime}\chi^{\prime}Q_{5},_{\chi}-\chi Q_{5},_{\phi}\right)\,,$ (17)
${\cal Q}$ $=$ $2\left[Q_{4}-2\chi Q_{4},_{\chi}+\chi\left(\frac{A^{\prime}}{2A}B\phi^{\prime}Q_{5},_{\chi}+Q_{5},_{\phi}\right)\right]\,,$ (18)
${\cal H}$ $=$ $2\left[Q_{4}-2\chi Q_{4},_{\chi}+\chi\left(\frac{C^{\prime}}{2C}B\phi^{\prime}Q_{5},_{\chi}+Q_{5},_{\phi}\right)\right]\,.$ (19)
To proceed, we introduce an auxiliary field and tortoise coordinate
$\Psi(r,t)$ $=\sqrt{\frac{BC\mathscr{H}^{2}Z}{A\mathscr{F}}}\left(\dot{h}_{1}-h^{\prime}_{0}+\frac{C^{\prime}h_{0}}{C}\right),$ (20)
$\mathrm{d}r_{*}$ $=\sqrt{\frac{Z^{2}}{AB}}\,\mathrm{d}r.$ (21)
It is noted that comparing against the formulas utilized in DeFelice:2011ka ; Ganguly:2017ort . Here we introduce an auxiliary function $Z$, whose form can be properly chosen to avoid the superluminal propagation of GWs, as we will see soon. After combining the transformations (20) and (21), the second order action (13) can be reduced as
$S^{(2)}=\frac{\ell(\ell+1)}{4(\ell-1)(\ell+2)}\int\mathrm{d}t\mathrm{d}r_{*}\left[\frac{\mathscr{F}}{\mathscr{G}Z^{2}}\left(\frac{\mathrm{d}\Psi}{\mathrm{d}t}\right)^{2}-\left(\frac{\mathrm{d}\Psi}{\mathrm{d}r_{*}}\right)^{2}-V(r)\Psi^{2}\right],$ (22)
where the effective potential $V(r)$ takes the form
$V(r)=\frac{(\ell-1)(\ell+2)A\mathscr{F}}{C\mathscr{H}Z^{2}}-\frac{C^{2}}{4C^{\prime}Z^{2}}\left(\frac{ABC^{\prime 2}}{C^{3}}\right)^{\prime}-\frac{C^{2}\mathscr{F}^{2}}{4\mathscr{F}^{\prime}}\left(\frac{AB\mathscr{F}^{\prime 2}}{C^{2}\mathscr{F}^{3}Z^{2}}\right)^{\prime}+\frac{\mathscr{F}^{2}}{Z^{\prime}}\left(\frac{ABZ^{\prime 2}}{4\mathscr{F}^{2}Z^{3}}\right)^{\prime}.$ (23)
It is obvious that such second order action is only well-defined for $\ell\geq 2$ but not for the modes $\ell=0$ and $\ell=1$ which are in fact not dynamical. $\ell=0$ is spherically symmetric and therefore obeys the Birkhoff theorem such that the perturbation solution is trivial. The second order action (22) is ill-defined for $\ell=1$ mode due to the overall term $(\ell-1)^{-1}$, so it should be specially treated. Fortunately, it was found that in GR $\ell=1$ mode corresponds to an infinitesimal shift of the position of the black hole Regge:1957td , while in general scalar-tensor theory $\ell=1$ mode could also correspond to a certain physical perturbation Kobayashi:2012kh , which will not be elaborately interpreted here because we will focus on the modes $\ell\geq 2$ in the current work.
By varying the action (22), we obtain the equation of motion
$\left(\frac{\mathrm{d}}{\mathrm{d}r_{*}^{2}}-\frac{\mathscr{F}}{\mathscr{G}Z^{2}}\frac{\mathrm{d}}{\mathrm{d}t^{2}}-V(r)\right)\Psi=0,$ (24)
which is the modified Regge-Wheeler equation for the black hole in the subclass of Horndeski theory with the action (2). The radial propagation speed of the odd perturbation can be defined by the coefficient of $\frac{\mathrm{d}}{\mathrm{d}t^{2}}$ as
$c^{2}_{r}=\frac{\mathscr{G}Z^{2}}{\mathscr{F}}.$ (25)
  • $Z=1$
    For $Z=1$, the previous derivation recovers those in DeFelice:2011ka ; Ganguly:2017ort . Especially, the radial propagation speed of the odd perturbation in the current Horndeski hairy background (11) is
    $c_{r}^{2}=\frac{\mathscr{G}}{\mathscr{F}}=\frac{Q_{4}-2\chi Q_{4,\chi}}{Q_{4}}=\left(1+\frac{3Q}{2r}\right)^{-1}.$ (26)
@@
    We can obviously see that a non-vanishing scalar charge $Q$ indeed causes the deviation of $c_{r}^{2}$ from the unity, i.e. the vacuum light speed in GR. Thus, the varying of parameter $Q$ may cause $c_{r}^{2}<0$, and Laplacian instability will occur. In order to avoid such instability, a condition for keeping a positive $c_{r}^{2}$ must be imposed at least at the event horizon, i.e.
@@
    which is consistent with the stability condition addressed in Kobayashi:2012kh ; Ganguly:2017ort . Moreover, to illustrate how the hair parameter affects the radial propagation speed of the odd perturbation in this case, we plot FIG.1 which shows that the propagation speed can be larger or smaller than the vacuum light speed depending on the hairy charge.
@@
  • $Z=\sqrt{\frac{\mathscr{F}}{\mathscr{G}}}=\sqrt{1+\frac{3Q}{2r}}$
    It is explicit that for $Z=\sqrt{\frac{\mathscr{F}}{\mathscr{G}}}$, the modified Regge-Wheeler equation (24) becomes a standard wavelike equation, and the propagation speed $c_{r}$ returns to unity, same as that in GR. Our analysis indicates that with proper auxiliary field and tortoise coordinate, the superluminal propagation in previous study Ganguly:2017ort can be mathematically cured.
Solving the equation (24) with proper boundary conditions allows us to extract the spectrum of the odd-parity perturbation and to further analyze the stability of the hairy black hole and other related properties. In principle, the transformations (20) and (21) are expected not to affect the dynamic of the perturbation. In order to check this, we shall solve the equations (24) for both $Z=1$ and $Z=\sqrt{1+\frac{3Q}{2r}}$ to extract the QNM spectrum, yielding the same results as shown in the next section.
@@ III Quasinormal mode
Recalling the Horndeski hairy black hole (11), the modified Regge-Wheeler equation (24) can be rewritten as
$\left[\frac{\mathrm{d}}{\mathrm{d}\bar{r}_{*}^{2}}-\left(1+\frac{3Q}{2r}\right)\frac{\mathrm{d}}{\mathrm{d}t^{2}}-V_{1}(r)\right]\Psi_{1}=0~{}~{}~{}~{}\text{with}~{}~{}~{}~{}\mathrm{d}\bar{r}_{*}=f^{-1}\mathrm{d}r,$ (28)
$V_{1}(r)=\frac{f^{2}\left(27Q^{2}+48Qr+32r^{2}\right)}{4r^{2}(3Q+2r)^{2}}+\left(1+\frac{3Q}{2r}\right)f\left[\frac{(\ell+2)(\ell-1)}{r^{2}}-\frac{f^{\prime}(3Q+4r)}{(3Q+2r)^{2}}\right]$ (29)
for $Z=1$, and
$\left[\frac{\mathrm{d}}{\mathrm{d}\tilde{r}_{*}^{2}}-\frac{\mathrm{d}}{\mathrm{d}t^{2}}-V_{2}(r)\right]\Psi_{2}=0~{}~{}~{}~{}\text{with}~{}~{}~{}~{}\mathrm{d}\tilde{r}_{*}=f^{-1}\left(1+\frac{3Q}{2r}\right)\mathrm{d}r,$ (30)
$V_{2}(r)=\frac{f^{2}\left(135Q^{2}+240Qr+128r^{2}\right)}{8r(3Q+2r)^{3}}+f\left[\frac{(\ell+2)(\ell-1)}{r^{2}}-\frac{f^{\prime}(9Q+8r)}{2(3Q+2r)^{2}}\right]$ (31)
for $Z=\sqrt{\frac{\mathscr{F}}{\mathscr{G}}}=\sqrt{1+\frac{3Q}{2r}}$. Both of them revert to the standard Regge-Wheeler potential Regge:1957td when the Horndeski hair vanishes. Since the explicit expression for the master equations are clearly established, we are now in a position to solve the equation and extract the dynamical properties for the odd-parity perturbation, which allows us to work in both the time and frequency domains.
@@ III.1 Preparation
Firstly, we will outline the method and boundary conditions in our numerical calculations to solve (28) in the time domain, and (28, 30) in the frequency domain.
@@
  • In the time domain, we shall work in tortoise coordinate grids and apply the finite difference method (Readers may refer to Zhu:2014sya ; Yang:2023lcm for more details) to discretize the radial coordinate $r=r(\bar{r}_{*})=r(j\Delta\bar{r}_{*})=r_{j}$, thus the perturbation and the potential are discretized as $\Psi(t,r)=\Psi(i\Delta t,j\Delta\bar{r}_{*})=\psi_{i,j}$ and $V(r)=V(j\Delta\bar{r}_{*})=V_{j}$. Then we use the central difference to discretize the derivative of the perturbation and rewrite the equation (28) into an iterative equation
    $\Psi_{i+1,j}=\frac{\Delta t^{2}\left(\Psi_{i,j-1}-2\Psi_{i,j}+\Psi_{i,j+1}\right)}{\left(1+\frac{3Q}{2r_{j}}\right)\Delta\bar{r}_{*}^{2}}-\Delta t^{2}V_{j}\Psi_{i,j}+2\Psi_{i,j}-\Psi_{i-1,j}.$ (32)
    To ensure reliable numerical results in solving partial differential equation, the Von Neumann stability condition, also known as Fourier stability analysis, should be analyzed to ensure that errors introduced during the computation do not grow uncontrollably as the calculation progresses. In this scenario, the Von Neumann stability condition usually requires $\frac{\Delta t}{\Delta\bar{r}_{*}}\leq 1$ which can make the numerical scheme stable Smith1985 , such that we set the intervals $\Delta\bar{r}_{*}=2\Delta t=0.1$. Moreover, we will choose the Gaussian packet $\Psi_{i<0,j}=0$, $\Psi_{0,j}=\exp[-\frac{(r_{j}-18)^{2}}{2}]$ as the initial data to evolve the iterative equation. The evolutionary profile of the perturbation typically includes an initial outburst, a damped oscillation, and a late-time tail, from which one can compare the damping time of the perturbation in varying black hole parameters and roughly predict the (in)stability of the background spacetime.
@@
  • In the frequency domain, we shall factor out the time dependency of the perturbations by separating $\Psi_{I}(t,r)=e^{-i\omega t}R_{I}(r)$ with $I=1,2$ denoting the fields in (28) and (30), which yield general eigenvalue equations
    $\left[\frac{\mathrm{d}}{\mathrm{d}\bar{r}_{*}^{2}}+\left(1+\frac{3Q}{2r}\right)\omega^{2}-V_{1}(r)\right]R_{1}(r)=0,~{}~{}~{}\text{for}~{}~{}Z=1,$ (33)
@@
    $\left[\frac{\mathrm{d}}{\mathrm{d}\tilde{r}_{*}^{2}}+\omega^{2}-V_{2}(r)\right]R_{2}(r)=0,~{}~{}~{}\text{for}~{}~{}Z=\sqrt{1+\frac{3Q}{2r}}.$ (34)
    Their corresponding boundary conditions for the QNMs solution are ingoing at the horizon
    $R_{1}(r\to r_{+})$ $\sim$ $e^{-i\sqrt{1+\frac{3Q}{2r_{+}}}\omega\bar{r}_{*}},$ (35)
    $R_{2}(r\to r_{+})$ $\sim$ $e^{-i\omega\tilde{r}_{*}},$ (36)
    and outgoing at the infinity
    $R_{1}(r\to\infty)\sim e^{i\omega\bar{r}_{*}},$ (37)
    $R_{2}(r\to\infty)\sim e^{i\omega\tilde{r}_{*}},$ (38)
@@
    respectively. Thus, once the black hole parameters (for instance, the mass $M$ and the hairy charge $Q$) are fixed, the eigenvalues could be solved as an infinite series of complex numbers, which are the so-called QNM frequencies. Similar to the normal modes, the real part of the QNMs represents the physical oscillation angular frequency, but they differ from the normal modes by having a non-vanishing imaginary part that represents the exponential damping of the perturbation in a dissipating system.
@@
    We shall employ the pseudospectral method and the matrix method to solve the master equation (33) with the bondary conditions (35) and (37), to firstly show the validity of the obtained QNMs. Then we apply the pseudospectral method to solve (34) with the boundary conditions (36) and (38), to verify the isospectrality between the two choices of $Z$. Since those algorithms are widely used in studies of black hole spectroscopy, here we shall skip the main steps, and readers may refer to Jansen:2017oag for the pseudospectral method and Lin:2019mmf for the matrix method respectively. It is noted that due to the $\log$ term in the metric function, a redefinition of the radial wave function is crucial when using both methods, as we have addressed in appendix B of our previous work Yang:2023lcm . Similar processes will be performed in the current numerical calculations.
@@
Following the above preparation, we shall show our numerical results of the eigenfrequencies of the QNMs and time evolution of perturbed field in next subsection. In our numeric, we find that as the hair parameter goes closer to the extremal case, the precision of our calculation becomes worse and even the code may break down. So, we shall focus on the parameter regime $-M\leq Q\leq M$, and for simplicity, we also rescale the Horndeski charge into the dimensionless $\tilde{Q}=Q/M$ in the calculations.
@@ III.2 Numerical results
The QNM frequencies with nodes $n=0,1$ for the selected hairy charges are listed in TABLE 1, where we ensure the accuracy of the results by evaluating their relative error. It is shown that for the $Z=1$ case, the pseudospectral and matrix methods achieve high accuracy in the nodes $n=0$ and $1$, and their results are in good agreement with each other, with a relative error on the order of $10^{-5}$ or even less. For the $Z=\sqrt{\mathscr{F}/\mathscr{G}}$ case, the QNM results obtained from the pseudospectral method are highly consistent with the results of the $Z=1$ case, with a relative error of less than $10^{-5}$. This demonstrates the isospectrality between the two choices of $Z$.
@@
$\ell=2$ $n=0$ $Z=1$ PS 0.747343 - 0.177925 i 0.679390 - 0.163833 i 0.617493 - 0.153284 i
M 0.747343 - 0.177925 i 0.679390 - 0.163833 i 0.617495 - 0.153284 i
$\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$
$Z=\sqrt{\mathscr{F}/\mathscr{G}}$ PS 0.747343 - 0.177925 i 0.679390 - 0.163833 i 0.617493 - 0.153284 i
$\mathscr{O}(10^{-11})$; $\mathscr{O}(10^{-11})$ $\mathscr{O}(10^{-9})$; $\mathscr{O}(10^{-8})$ $\mathscr{O}(10^{-8})$; $\mathscr{O}(10^{-7})$
$n=1$ $Z=1$ PS 0.693422 - 0.547830 i 0.623772 - 0.504366 i 0.573985 - 0.481447 i
M 0.693420 - 0.547829 i 0.623768 - 0.504368 i 0.574023 - 0.481421 i
$\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-5})$; $\mathscr{O}(10^{-5})$
$Z=\sqrt{\mathscr{F}/\mathscr{G}}$ PS 0.693422 - 0.547830 i 0.623772 - 0.504366 i 0.573993 - 0.481448 i
$\mathscr{O}(10^{-10})$; $\mathscr{O}(10^{-10})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-8})$ $\mathscr{O}(10^{-5})$; $\mathscr{O}(10^{-7})$
$\ell=3$ $n=0$ $Z=1$ PS 1.198887 - 0.185406 i 1.078947 - 0.167580 i 0.955588 - 0.150350 i
M 1.198887 - 0.185406 i 1.078947 - 0.167580 i 0.955587 - 0.150349 i
$\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$
$Z=\sqrt{\mathscr{F}/\mathscr{G}}$ PS 1.198887 - 0.185406 i 1.078947 - 0.167580 i 0.955588 - 0.150350 i
$\mathscr{O}(10^{-13})$;$\mathscr{O}(10^{-11})$ $\mathscr{O}(10^{-9})$; $\mathscr{O}(10^{-8})$ $\mathscr{O}(10^{-8})$; $\mathscr{O}(10^{-8})$
$n=1$ $Z=1$ PS 1.165288 - 0.562596 i 1.046267 - 0.508376 i 0.929446 - 0.458644 i
M 1.165288 - 0.562596 i 1.046269 - 0.508377 i 0.929442 - 0.458645 i
$\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$
$Z=\sqrt{\mathscr{F}/\mathscr{G}}$ PS 1.165288 - 0.562596 i 1.046267 - 0.508376 i 0.929442 - 0.458641 i
$\mathscr{O}(10^{-11})$; $\mathscr{O}(10^{-11})$ $\mathscr{O}(10^{-7})$; $\mathscr{O}(10^{-7})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$
$\ell=4$ $n=0$ $Z=1$ PS 1.618357 - 0.188328 i 1.451873 - 0.169314 i 1.274651 - 0.149665 i
M 1.618357 - 0.188328 i 1.451873 - 0.169314 i 1.274651 - 0.149666 i
$\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$
$Z=\sqrt{\mathscr{F}/\mathscr{G}}$ PS 1.618357 - 0.188328 i 1.451873 - 0.169314 i 1.274651 - 0.149665 i
$\mathscr{O}(10^{-13})$; $\mathscr{O}(10^{-12})$ $\mathscr{O}(10^{-9})$; $\mathscr{O}(10^{-9})$ $\mathscr{O}(10^{-9})$; $\mathscr{O}(10^{-8})$
$n=1$ $Z=1$ PS 1.593263 - 0.568669 i 1.427902 - 0.511200 i 1.255580 - 0.452830 i
M 1.593263 - 0.568669 i 1.427904 - 0.511200 i 1.255583 - 0.452831 i
$\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$ $\mathscr{O}(10^{-6})$; $\mathscr{O}(10^{-6})$
$Z=\sqrt{\mathscr{F}/\mathscr{G}}$ PS 1.593263 - 0.568669 i 1.427902 - 0.511201 i 1.255579 - 0.452831 i
$\mathscr{O}(10^{-11})$; $\mathscr{O}(10^{-11})$ $\mathscr{O}(10^{-8})$; $\mathscr{O}(10^{-7})$ $\mathscr{O}(10^{-7})$; $\mathscr{O}(10^{-6})$
Table 1: The QNMs solved from the equations of the odd perturbation (24) of the non-rotating Horndeski hairy black hole (10), for $\ell=2,3,4$ and $n=0,1$. For the $Z=1$ case, the QNM data are obtained by the pseudospectral (abbreviation as PS) and matrix (abbreviation as M) methods, and their relative errors are evaluated as $\Delta_{k}=\mid\frac{\text{M}(\omega_{k})-\text{PS}(\omega_{k})}{\text{PS}(\omega_{k})}\mid$ with $k=\text{Re},\text{Im}$. For the $Z=\sqrt{\mathscr{F}/\mathscr{G}}$ case, the data are determined by the pseudospectral method, and their errors are defined as the deviation of QNMs frequencies between two choices of $Z$, i.e. $\Delta_{k}=\mid\frac{\text{PS}(\omega^{\prime}_{k})-\text{PS}(\omega_{k})}{\text{PS}(\omega_{k})}\mid$. In the cells, we give the corresponding errors for real and imaginary of the frequencies, repectively.